1) At 100 oC, Keq = 1.5E8 for the reaction:

CO(g) + Cl2(g) COCl2(g)

Using appropriate approximation, calculate the partial pressure of CO at 100 oC at equilibrium in a chamber that initially contains COCl2 at a pressure of 0.293 bar.
Answer #1

first construct the ICE table and then substitute the values in the equation and then calculate the equilibrium values as follows
El cock 0.293 Elli al coll tx tx -Xe 009 6.292-x) k= [cocka ] [co] [dz] -) 1.5xlof - (0.299-X) -> (1.57109) x2 = (0.293-6) =>